# APEX-MR: Multi-Robot Asynchronous Planning and Execution for Cooperative Assembly

### v1

@@ IV-D TPG Execution
Executing a motion plan on robot arms often requires a position controller for movement and other specific controllers for manipulation skills. These controllers may have delays or uncertainties that affect the real-robot execution time. The TPG formulation provides an easy way to execute a multi-robot plan. Here, we present one centralized mechanism to coordinate multiple robot arms.
As shown in Fig. 4, the TPG is hosted on a central server that communicates with each robot’s execution thread. Each pose node in the TPG corresponds to an action that moves the robot’s position to the node’s configuration. Each skill node corresponds to an action that executes the predefined robot skill. An action can be safely executed if there are no incoming edges from any nodes that are not executed.
If an action is safe to execute based on the TPG, the central server sends it to the robot’s action queue. Each robot maintains its own controller-sensing loop and actuates the robot according to upcoming commands and its state estimation. The state estimation is also shared with the TPG, which then updates the TPG when a node is being executed or completed. Newly completed nodes may also enqueue new nodes if their outgoing edges were previously preventing unsafe actions. During execution, TPG can be interpreted as a control law that maintains the safe scheduling of individual robot actions.
@@ V Results
To evaluate the performance, we apply the proposed APEX-MR to bimanual LEGO assembly tasks. Given a customized LEGO design as shown in LABEL:fig:eval_lego, APEX-MR coordinates the robots to construct the desired structure as shown in LABEL:fig:lego_assembly_structures using available LEGO bricks. We deliberate the inputs (introduced in Sec. III) to APEX-MR below.
@@
Environment Setup LABEL:fig:sim_env_setup and LABEL:fig:env_setup illustrate the simulation environment and the real setup, which includes two Yaskawa GP4 robots. Following the task convention in [30], we consider building LEGO structures on a baseplate, which is calibrated [^1] We calibrate the transformation from the robots to the baseplate by teleoperating the robot to touch the plate. Note that we only measure the translation ($X$, $Y$, $Z$) and yaw angle while assuming no roll and pitch offsets. and placed in between the two robots, using commercial standard LEGO bricks initially stored on the baseplate. Each robot is equipped with an ATI Gamma force-torque sensor (FTS) and the end-of-arm tool (EOAT) is mounted on the FTS. The simulation consists of the entire workspace, which includes the robots, FTS, EOAT, LEGOs, nearby workstations, etc. The complete digital environment provides rich and accurate information for APEX-MR to safely coordinate the robot collaboration.
@@
Assembly Plan Given a LEGO structure, we employ the physics-aware assembly planning in [28] with customized LEGO physics reasoning [29] to generate a physically valid assembly sequence. Specifically, a physically valid assembly sequence enforces that for each step after assembling a brick, the structure is stable and does not collapse. Note that the definition of a physically valid assembly sequence can be different for other cooperative assembly tasks. For other applications, the assembly sequence can be obtained via planners, e.g., [58], and the proposed APEX-MR is also applicable downstream.
@@
Manipulation Skills Manipulating LEGO bricks is a non-trivial contact-rich manipulation problem beyond simple pick and stack. A robot EOAT and manipulation policy (i.e., insert-and-twist) was presented in [30], which enables a robot to manipulate commercial standard LEGO bricks, i.e., pick, and place-down in Fig. 1. However, a robot can only use it to manipulate a LEGO brick from its top, which limits the system from constructing complex structures. To enhance the system capability, we present a new LEGO tool (LT-V2), as shown in LABEL:fig:eoat. In particular, LT-V2 has LEGO studs added to the side of the tooltip. The new design enables the robot to manipulate a brick from its bottom as shown in Fig. 1, i.e., handover and place-up. With LT-V2, we define the manipulation skills as shown in Fig. 1, including 1) goal reaching with force feedback (i.e., support-bottom and support-top), and 2) learned force policy (i.e., pick, place-down, place-up, handover). For each skill, we generate a maximum of one feasible LEGO grasp pose and one support pose if necessary to use in task planning.
@@
Figure 7: Normalized makespan and wait time of APEX-MR (Sequential) versus the synchronized planner across example evaluation environment. The results are normalized by the makespan of the sequential motion plan before TPG and averaged over 4 random seeds. The unnormalized makespan and wait time are labeled for each entry and the dashed horizontal line corresponds to the makespan of the sequential motion plan before TPG.
@@
Figure 8: Breakdown of wall clock time of APEX-MR (Sequential) and the synchronized planner baseline across evaluation environments. The results are normalized by the total of APEX-MR planning time for each task (excluding shortcutting) and averaged over 4 random seeds. The running sum of unnormalized wall clock time is labeled for each component.
@@
Implementation We implement the TPG algorithm and manipulation skills in C++ with ROS-Noetic and MoveIt [7]. The ILP in task planning is solved with the pulp Python package. The RRT-Connect motion planning uses MoveIt’s OMPL [57] plug-in. Trajectories are discretized using $\Delta t=0.05$ when the maximum $L_{1}$ joint velocity is 1 rad/s. $\Delta t$ is adjusted linearly based on the max velocity to ensure the same density. All simulation experiments are conducted on an AMD 7840HS laptop. Algo. 1 parallelizes collision checking with 16 threads.
@@
Experiment Objective While APEX-MR itself is a full pipeline for multi-robot tasks and motion planning, the key innovation that enables asynchronous innovation is the TPG execution framework. Thus, we are interested in the following questions when evaluating APEX-MR:
@@
  • (Q3) How well does APEX-MR perform in physical LEGO assembly and can it safely execute planned trajectories despite uncertainties?
Q1 and Q2 will be examined closely in simulation, whereas Q3 will be the focus of our real robot experiments.
@@ V-A Simulation Performance
Metrics We use the execution makespan and wait time as our evaluation metrics for plan quality. Since our output is a TPG, we first roll out the asynchronous trajectory from the TPG, assuming no controller delay. The rollout trajectory converts each pose node back to a configuration. Actions in the skill nodes are executed based on the reference trajectory, and force feedback is switched off in the simulation. The timestamp for each configuration in the rollout trajectory is the earliest possible time to reach this node based on incoming type-2 edges. Execution makespan is the maximum time taken among all robots to execute a trajectory, i.e., $\max_{i=1}^{N}t^{i}_{N^{i}_{end}}$. Wait time is defined as the total amount of time any robot spends waiting in the rollout asynchronous trajectory from TPG, or the original sequential or synchronous trajectory. In particular, we are interested in whether TPG processing can successfully reduce wait time when initialized with a sequential task and motion plan.
@@
Baseline We also design a baseline for a synchronized task and motion planning as a comparison to APEX-MR. Although APEX-MR uses a sequential planner for simplicity and efficiency, our TPG can also improve synchronized motion plans which is a common approach in MR-TAMP. Thus, we evaluate the performance improvement of TPG on synchronized plan as well. In this synchronous planner, robots can execute tasks in parallel but must wait for all robots to finish their current task before proceeding to the next set of tasks. We use an algorithm to convert the sequential task plan from APEX-MR to a synchronous task plan, as shown by Fig. 15 in the Appendix. The main idea is to execute the sequential task plan in parallel if executing the next task does not violate inter-robot task dependencies or block tasks scheduled at an earlier time. This process is similar to building a TPG on tasks instead of motions for parallelization. For every robot $i$ and its task $\mathcal{T}_{m}^{i}$, the algorithm checks if the intermediate goal pose $C^{i}_{N_{end,m}^{i}}$ collides with any other intermediate goal pose of robot $i^{\prime}$ and task $\mathcal{T}_{m^{\prime}}^{i^{\prime}}$ that satisfies $m^{\prime}<m$. If there exists a collision, then robot $i$ must wait for robot $i^{\prime}$ to complete task $\mathcal{T}_{m^{\prime}}^{i^{\prime}}$ before starting task $\mathcal{T}_{m}^{i}$. A synchronous task graph can then be generated by combining these calculated dependencies, with existing inter-robot dependencies. Then, composite RRT-Connect is used as the multi-robot motion planner. Synchronous trajectories for tasks executing in parallel are generated by planning all degrees of freedom as a single robot.
@@
Performance Fig. 7 shows the quality of the solution of APEX-MR on a variety of tasks. First, the TPG post-processing and applying and shortcut, as described in Sec. IV-C, significantly reduces makespan by 48% and wait time by 85% on average, compared to the initial sequential motion plan at the horizontal dashed line. Compared to the synchronized motion plan, our asynchronous plans from APEX-MR are consistently shorter and have less wait time. When applied to the synchronous plan, TPG also significantly reduces the makespan by 36% and wait time by 77% on average. Note that the post-processed sequential plan from APEX-MR still slightly outperforms the post-processed synchronized motion plan by 3% in terms of makespan. This is due to the path produced by a multi-robot motion plan being suboptimal compared to sequential motion planning. Still, the wait time for the synchronized plan after TPG post-processing is minimal.
@@
Runtime Fig. 8 and Table I in the appendix shows the wall clock time for APEX-MR. On average, the TPG construction time is always lower than the task motion planning time except for the Chair, which has a very long assembly sequence. On the other hand, running a synchronized planner can be much more expensive than the simple sequential planner used in APEX-MR, which requires more careful coordination in task planning, and multi-robot motion planning with more degrees of freedom. By combining a simple sequential task and motion planner with TPG post-processing, APEX-MR produces higher-quality multi-robot plans with 26% lower computational overhead on average than a synchronized multi-robot task and motion planner alone.
@@ V-B System Deployment
Note that the key difference between assembling in real and in simulation is with and without force feedback. Integrating force feedback into the manipulation skills improves the system robustness, but also brings uncertainty with respect to execution time. In particular, we are interested in whether the TPG execution framework ensures safe execution and avoids collisions despite uncertainties from manipulation skills. Fig. 14 depicts the distribution of execution times of six manipulation skills, i.e., pick, place-down, place-up, handover, support-bottom, and support-top (see Fig. 1 for illustrations). All of these skills are designed to execute with force feedback to ensure proper contact between the EOAT and the object being manipulated and minimize the effect of imperfect calibration or tilted structure. As a consequence of these mm-level adjustments with force feedback, the execution time can vary as much as 2 seconds, or as much as 23% from the median. Nevertheless, our TPG execution framework can reliably adjust the robot schedule if any delay could cause a collision or require another robot to wait longer until a skill is completed.
@@ VI Limitations and Future Work
Offline Computation Currently, both the TPG processing and motion planning in APEX-MR are performed offline before real execution. This can be a drawback in real assembly tasks, where new tasks are continuously assigned once the robots finish existing assembly on an assembly line. Another limitation of offline computation is that the APEX-MR cannot adapt to changes in the collision environment or assembly steps easily. A principled framework to address these lifelong planning techniques is to use a windowed multi-robot planner and only convert the first $n$ part of the robot plan to a TPG, similar to how Varambally et al. [59] address automated warehousing. Taking a reduced-horizon approach will significantly reduce planning time and allow plans to be continuously updated concurrently with execution.
@@
Planning for Robot Dynamics While APEX-MR are reliable and safe on real robots, APEX-MR requires a good position or force controller for the robot because the generated plan does not consider robot dynamics, such as acceleration and jerk constraints. This is challenging because the planner must generate continuous velocity and acceleration profiles while also avoiding inter-robot collisions at all times. We plan to incorporate dynamics as part of the TPG post-processing step, such as solving a linear program on top of TPG-imposed constraints and velocity constraints, as suggested in Hönig et al. [20]. Another interesting problem with robot arms is the speed constraints may be imposed on the task space, due to the attached object at the end-effector or even closed kinematic chains formed by concurrent manipulation [60].
@@
Manipulation Policy While the proposed APEX-MR enables efficient and safe dual-arm cooperative LEGO assembly, the current system assembles each object based on pre-defined skills. With the additional force feedback, each single operation can be performed robustly. However, due to the passive connection nature of LEGO structures, i.e., established connections could be gradually loosened and the structure can be tilted or even collapse due to subsequent operations, the long-horizon assembly could still fail. Therefore, the dual-arm system, at its current stage, is not robust enough to construct large-scale LEGO structures that have multiple fragile overhanging geometries. To further improve the robustness from a system perspective, we aim to investigate methods for failure detection [4] and recovery, e.g., reinforce the connections that are loosened due to later operations. Failure detection and recovery can also be integrated as part of an online replanning framework that dynamically reschedules the robot’s tasks if a failure occurs and intervention becomes necessary.
@@
Other Cooperative Assemblies Building on the APEX-MR pipeline, we plan to extend its application to a broader range of cooperative tasks, such as the NIST Box Assembly [40], and other industrial assembly scenarios. By doing so, we aim to address the unique challenges posed by real-world manufacturing environments, gaining deeper insights into how cooperative systems can be optimized for complex, large-scale production tasks. While the discussion in this paper is based on LEGO assembly, the components in APEX-MR, especially TPG post-processing, can be applied to other multi-robot assembly tasks. A concrete step would be to investigate how to integrate manipulation policies that are more complex than those used for LEGO assembly, e.g., diffusion policy [6], to the TPG framework.

### v2

@@ IV-D TPG Execution
Executing a motion plan on robot arms often requires a position controller for movement and other specific controllers for manipulation skills. These controllers may have delays or uncertainties that affect the real-robot execution time. The TPG formulation provides an easy way to execute a multi-robot plan. Here, we present a semi-centralized mechanism to coordinate multiple robot arms.
As shown in Fig. 4, the TPG is hosted on a central server that communicates with each robot’s execution thread. Each pose node in the TPG corresponds to an action that moves the robot’s position to the node’s configuration. Each skill node corresponds to an action that executes the predefined robot skill. An action can be executed safely if there are no incoming edges from any nodes that are not executed. Although each TPG node is associated with a timestamp, it is only used for TPG construction and is ignored during execution.
If an action is safe to execute based on the TPG, the central server sends it to the robot’s action queue. Each robot maintains its own controller-sensing loop and actuates the robot according to upcoming commands and its state estimation. The state estimation is also shared with the TPG, which then updates the TPG when a node is being executed or completed. Newly completed nodes can enable the central server to enqueue new nodes if their outgoing edges were previously preventing unsafe actions. During execution, TPG can be interpreted as a control law that maintains the safe scheduling of individual robot actions.
@@ V Results
To evaluate performance, we apply the proposed APEX-MR to bimanual LEGO assembly tasks. Given a customized LEGO design as shown in LABEL:fig:eval_lego, APEX-MR coordinates the robots to construct the desired structure as shown in LABEL:fig:lego_assembly_structures using available LEGO bricks. We deliberate the inputs (introduced in Sec. III) to APEX-MR below.
@@
Environment Setup LABEL:fig:sim_env_setup and LABEL:fig:env_setup illustrate the simulation environment and the real setup, which includes two Yaskawa GP4 robots. Following the task convention in [29], we consider building LEGO structures on a baseplate, which is calibrated [^3] We calibrate the transformation from the robots to the baseplate by teleoperating the robot to touch the plate. Note that we only measure the translation ($X$, $Y$, $Z$) and yaw angle while assuming no roll and pitch offsets. and placed between the two robots, using commercial standard LEGO bricks initially stored on the baseplate. Each robot is equipped with an ATI Gamma force-torque sensor (FTS), and the end-of-arm tool (EOAT) is mounted on the FTS. The simulation consists of the entire workspace, which includes robots, FTS, EOAT, LEGOs, nearby workstations, etc. The complete digital environment provides rich and accurate information for APEX-MR to safely coordinate robot collaboration.
@@
Assembly Plan Given a LEGO structure, we employ the physics-aware assembly planning in [30] with customized LEGO physics reasoning [28] to generate a physically valid assembly sequence. Specifically, a physically valid assembly sequence enforces that for each step after assembling a brick, the structure is stable and does not collapse. More details on generating the physically valid assembly sequence are provided in Sec. -A. Note that the definition of a physically valid assembly sequence can be different for other cooperative assembly tasks. For other applications, the assembly sequence can be obtained via planners, e.g., [58], and the proposed APEX-MR is also applicable downstream.
@@
Manipulation Skills Manipulating LEGO bricks is a non-trivial contact-rich manipulation problem beyond simple pick and stack. A robot EOAT and manipulation policy (i.e., insert-and-twist) were presented in [29], which enable a robot to manipulate commercial standard LEGO bricks, i.e., pick, and place-down in Fig. 1. However, a robot can only use it to manipulate a LEGO brick from its top, which limits the system from constructing complex structures. To enhance the system capability, we present a new LEGO tool (LT-V2), as shown in LABEL:fig:eoat. In particular, LT-V2 has LEGO studs added to the side of the tooltip. The new design enables the robot to manipulate a brick from its bottom as shown in Fig. 1, i.e., handover and place-up. With LT-V2, we define the manipulation skills as shown in Fig. 1, including 1) goal reaching with force feedback (i.e., support-bottom and support-top), and 2) learned force policy (i.e., pick, place-down, place-up, handover). More details on LEGO manipulation skills are discussed in Sec. -C.
@@
Figure 7: Normalized makespan and wait time of APEX-MR (Sequential) versus the synchronized planner across example evaluation environment. The results are normalized by the makespan of the sequential motion plan before TPG and averaged over 4 random seeds. The unnormalized makespan and wait time in seconds are labeled for each entry, and the dashed horizontal line corresponds to the makespan of the sequential motion plan before TPG. Shortcut refers to the anytime shortcutting algorithm on TPG in Sec. IV-C.
@@
Figure 8: Breakdown of wall clock time of APEX-MR (Sequential) and the synchronized planner baseline across evaluation environments. The results are normalized by the total of APEX-MR planning time for each task (excluding shortcutting) and averaged over 4 random seeds. The running sum of unnormalized wall clock time in seconds is labeled for each component.
@@
Implementation We implement the TPG algorithm and manipulation skills in C++ with ROS-Noetic and MoveIt [7]. The ILP in task planning is solved with the pulp Python package. The RRT-Connect motion planning uses MoveIt’s OMPL [57] plug-in. Paths are discretized using $\Delta t=0.05$ seconds when the maximum $L_{1}$ joint velocity is 1 rad/s. $\Delta t$ is adjusted linearly based on the maximum velocity to ensure the same density. All simulation experiments are conducted on an AMD 7840HS laptop. Algo. 1 parallelizes collision checking with 16 threads.
@@
Experiment Objective While APEX-MR itself is a full pipeline for multi-robot tasks and motion planning, the key innovation that enables asynchronous collaboration is the TPG execution framework. Thus, we are interested in the following questions when evaluating APEX-MR:
@@
  • (Q3) How well does APEX-MR perform in physical LEGO assembly, and can it safely execute planned paths despite uncertainties?
Q1 and Q2 will be closely examined in simulation, whereas Q3 will be the focus of our real robot experiments.
@@ V-A Simulation Performance
Metrics We use execution makespan and wait time as our evaluation metrics for plan quality. Since our output is a TPG, we first roll out the asynchronous path from the TPG, assuming no controller delay. The rollout path converts each pose node back to a configuration. Actions in the skill nodes are executed based on the reference path, and force feedback is switched off in the simulation. The timestamp for each configuration in the rollout path is the earliest possible time to reach this node based on incoming type-2 edges. Execution makespan is the maximum time taken among all robots to execute their rollout paths, i.e., $\max_{i=1}^{N}t^{i}_{N^{i}_{end}}$. Wait time is defined as the total amount of time any robot spends waiting in the rollout asynchronous path from TPG, or the original sequential or synchronous path. In particular, we are interested in whether TPG processing can successfully reduce wait time when initialized with a sequential task and motion plan.
@@
Baseline We also design a baseline for synchronized task and motion planning as a comparison to APEX-MR. Although APEX-MR uses a sequential planner for simplicity and efficiency, our TPG can also improve synchronized motion plans, common in MR-TAMP. Thus, we evaluate the performance improvement of TPG on synchronized plans. In this synchronous planner, robots can execute tasks in parallel but must wait for all robots to finish their current task before proceeding to the next set of tasks. We use an algorithm to convert the sequential task plan from APEX-MR to a synchronous task plan, as shown by Fig. 17 in Sec. -E. The main idea is to execute the sequential task plan in parallel if executing the next task does not violate inter-robot task dependencies or block tasks scheduled at an earlier time. This process is similar to building a TPG on tasks instead of motions for parallelization. For every robot $i$ and its task $\mathcal{T}_{m}^{i}$, the algorithm checks if the intermediate goal pose $C^{i}_{N_{end,m}^{i}}$ collides with any other intermediate goal pose of robot $i^{\prime}$ and task $\mathcal{T}_{m^{\prime}}^{i^{\prime}}$ that satisfies $m^{\prime}<m$. If there exists a collision, then robot $i$ must wait for robot $i^{\prime}$ to complete task $\mathcal{T}_{m^{\prime}}^{i^{\prime}}$ before starting task $\mathcal{T}_{m}^{i}$. A synchronous task graph can then be generated by combining these calculated dependencies with existing inter-robot dependencies. Then, composite RRT-Connect is used as the multi-robot motion planner. Synchronous paths for tasks executing in parallel are generated by planning all degrees of freedom as a single robot.
@@
Performance Fig. 7 shows the quality of the solution of APEX-MR on a variety of tasks. First, the TPG post-processing and applying and shortcut, as described in Sec. IV-C, significantly reduces makespan by 48% and wait time by 85% on average, compared to the initial sequential motion plan on the horizontal dashed line. Compared to the synchronized motion plan, our asynchronous plans from APEX-MR are consistently shorter and have lower wait time. When applied to the synchronous plan, TPG also significantly reduces the makespan by 36% and wait time by 77% on average. Note that the post-processed sequential plan from APEX-MR still slightly outperforms the post-processed synchronized motion plan by 3% in terms of makespan. This is due to the path produced by a multi-robot motion plan being suboptimal compared to sequential motion planning. Still, the wait time for the synchronized plan after TPG post-processing is minimal.
@@
Runtime Fig. 8 and Table II in Sec. -F shows the wall clock time for APEX-MR. On average, the TPG construction time is always lower than the task motion planning time except for the Chair, which has a very long assembly sequence. On the other hand, running a synchronized planner can be much more expensive than the simple sequential planner used in APEX-MR, which requires more careful coordination in task planning and multi-robot motion planning with more degrees of freedom. By combining a simple sequential task and motion planner with TPG post-processing, APEX-MR produces higher-quality multi-robot plans with 26% lower computational overhead on average than a synchronized multi-robot task and motion planner alone. One concern is that the number of collision checks when building TPG scales quadratically with the number of robots. Thus, we perform additional experiments in Sec. -D with three and four robot arms to show that the runtime of TPG construction remains reasonable.
@@ V-B System Deployment
Note that the key difference between assembling in real and in simulation is with and without force feedback. Integrating force feedback into the manipulation skills improves the system robustness, but also brings uncertainty with respect to execution time. In particular, we are interested in whether the TPG execution framework ensures safe execution and avoids collisions despite uncertainties of manipulation skills. Fig. 14 depicts the distribution of execution times of six manipulation skills, i.e., pick, place-down, place-up, handover, support-bottom, and support-top (see Fig. 1 for illustrations). All of these skills are designed to execute with force feedback to ensure proper contact between the EOAT and the object being manipulated and minimize the effect of imperfect calibration or a tilted structure. As a consequence of these mm-level adjustments with force feedback, the execution time can vary as much as 2 seconds, or as much as 23% from the median. Nevertheless, our TPG execution framework can reliably adjust the robot schedule if any delay could cause a collision or require another robot to wait longer until a skill is completed.
@@ VI Limitations and Future Work
Offline Computation Currently, both the TPG processing and motion planning in APEX-MR are performed offline before real execution. This can be a drawback in real assembly tasks, where new tasks are continuously assigned once the robots finish an existing assembly on an assembly line. Another limitation of offline computation is that the APEX-MR cannot easily adapt to changes in the collision environment or assembly steps. A principled framework to address these lifelong planning techniques is to use a windowed multi-robot planner and only convert the first $n$ steps of the robot plan to a TPG, similar to how Varambally et al. [59] address mobile robot coordination in automated warehousing. Taking a reduced-horizon approach will significantly reduce planning time and allow plans to be continuously updated concurrently with execution.
@@
Planning for Robot Dynamics While APEX-MR are reliable and safe on real robots, APEX-MR requires a good position or force controller for the robot because the generated plan does not consider robot dynamics, such as acceleration and jerk constraints. This is challenging because the planner must generate continuous velocity and acceleration profiles while also avoiding inter-robot collisions at all times. We plan to incorporate dynamics as part of the TPG post-processing step, such as solving a linear program on top of TPG-imposed constraints and velocity constraints, as suggested in Hoenig et al. [19]. Another interesting problem with robot arms is that speed constraints may be imposed on the task space, due to the attached object at the end-effector or even closed kinematics chains formed by concurrent manipulation [60].
@@
Manipulation Policy While the proposed APEX-MR enables efficient and safe dual-arm cooperative LEGO assembly, the current system assembles each object based on predefined skills. With the additional force feedback, each single operation can be performed robustly. However, due to the passive connection nature of LEGO structures, i.e., established connections could be gradually loosened and the structure can be tilted or even collapse due to subsequent operations, the long-horizon assembly could still fail. Therefore, the dual-arm system, at its current stage, is not robust enough to construct large-scale LEGO structures that have multiple fragile overhanging geometries. To further improve the robustness from a system perspective, our aim is to investigate methods for failure detection [4] and recovery, e.g., reinforcing the connections that are loosened due to later operations. Failure detection and recovery can also be integrated as part of an online replanning framework that dynamically reschedules the robot’s tasks if a failure occurs and intervention becomes necessary.
@@
Other Cooperative Assemblies Building on the APEX-MR pipeline, we plan to extend its application to a broader range of cooperative tasks, such as the NIST Box Assembly [40], and other industrial assembly scenarios. In doing so, we aim to address the unique challenges posed by real-world manufacturing environments, gaining deeper insights into how cooperative systems can be optimized for complex, large-scale production tasks. While the discussion in this paper is based on LEGO assembly, the components in APEX-MR, especially TPG post-processing, can be applied to other multi-robot assembly tasks. A concrete step would be to investigate how to integrate manipulation policies that are more complex than those used for LEGO assembly, e.g., diffusion policy [6], to the TPG framework.